OpenStudy (anonymous):

complete the square (1/4x^2)+6x+7

OpenStudy (anonymous):

That's \(\dfrac{1}{4}x^2\), and not \(\dfrac{1}{4x^2}\), correct?

OpenStudy (anonymous):

\[\begin{align*}\frac{1}{4}x^2+6x+7&=\frac{1}{4}\left(x^2+24x+\frac{7}{4}\right)\\ &=\frac{1}{4}\left(x^2+24x+144-144+\frac{7}{4}\right)\\ &=\frac{1}{4}\left((x+12)^2+\cdots\right) \end{align*}\]